# Gödel Agent: A Self-Referential Framework for Agents Recursively Self-Improvement

### v2

@@ Abstract
The rapid advancement of large language models (LLMs) has significantly enhanced the capabilities of AI-driven agents across various tasks. However, existing agentic systems, whether based on fixed pipeline algorithms or pre-defined meta-learning frameworks, cannot search the whole agent design space due to the restriction of human-designed components, and thus might miss the globally optimal agent design. In this paper, we introduce Gödel Agent, a self-evolving framework inspired by the Gödel machine, enabling agents to recursively improve themselves without relying on predefined routines or fixed optimization algorithms. Gödel Agent leverages LLMs to dynamically modify its own logic and behavior, guided solely by high-level objectives through prompting. Experimental results on multiple domains including coding, science, and math demonstrate that implementation of Gödel Agent can achieve continuous self-improvement, surpassing manually crafted agents in performance, efficiency, and generalizability [^1] Our code is released at https://github.com/Arvid-pku/Godel_Agent.
@@ 1 Introduction
As large language models (LLMs) such as GPT-4 (OpenAI et al., 2024) and LLaMA3(Dubey et al., 2024) demonstrate increasingly strong reasoning and planning capabilities, LLM-driven agentic systems have achieved remarkable performance in a wide range of tasks (Wang et al., 2024a). Substantial effort has been invested in manually designing sophisticated agentic systems using human priors in different application areas. Recently, there has been a significant interest in creating self-evolving agents with minimal human effort, which not only greatly reduces human labor but also produces better solutions by incorporating environmental feedback. Given that human effort can only cover a small search space of agent design, it is reasonable to expect that a self-evolving agent with the freedom to explore the full design space has the potential to produce the global optimal solution.
@@
There is a large body of work proposing agents capable of self-refinement. However, there are inevitably some human priors involved in these agent designs. Some agents are designed to iterate over a fixed routine consisting of a list of fixed modules, while some of the modules are capable of taking self- or environment feedback to refine their actions  (Shinn et al., 2024; Chen et al., 2023b; Qu et al., 2024; Yao et al., 2023). This type of agent, referred to as Hand-Designed Agent, is depicted as having the lowest degree of freedom in Figure 1. More automated agents have been designed to be able to update their routines or modules in some pre-defined meta-learning routine, for example, natural language gradients (Zhou et al., 2024), meta agent (Hu et al., 2024), or creating and collecting demonstrations  (Khattab et al., 2023). This type of agent, known as Meta-Learning Optimized Agents, is depicted as having the middle degree of freedom in Figure 1.
It is evident that both types of agents above are inherently constrained by human priors and one intuitional method to further increase the freedom of self-improvement is to design a meta-meta-learning algorithm, to learn the meta-learning algorithm. However, there is always a higher-level meta-learning algorithm that can be manually designed to learn the current-level meta-learning method, creating a never-ending hierarchy of meta-learning.
In this paper, we propose Gödel Agent to eliminate the human design prior, which is an automated LLM agent that can freely decide its own routine, modules, and even the way to update them. It is inspired by the self-referential Gödel machine (Schmidhuber, 2003), which was originally proposed to solve formal proof problems and was proven to be able to find the global optimal solutions. Self-reference means the property of a system that can analyze and modify its own code, including the parts responsible for the analysis and modification processes (Astrachan, 1994). Therefore, it can achieve what’s known as ”recursive self-improvement”, where it iteratively updates itself to become more efficient and effective at achieving its predefined goals. In this case, Gödel Agent can analyze and modify its own code, including the code for analyzing and modifying itself, and thus can search the full agent design space, which is depicted as having the highest degree of freedom in Figure 1. Gödel Agent can theoretically make increasingly better modifications over time through recursively self-update (Yampolskiy, 2015; Wang, 2018).
@@
Figure 1: Comparison of three agent paradigms. Hand-designed agents rely on human expertise which are limited in scope and labor-intensive. Meta-learning optimized agents are constrained by a fixed meta-learning algorithm, restricting their search space and optimization potential. In contrast, self-referential agent (Gödel Agent) can recursively improve itself without any limitation. Note that the input to Gödel Agent is itself, allowing it to modify itself and output a new version of itself.
@@
In this paper, we choose to implement it by letting it manipulate its own runtime memory, i.e., the agent is able to retrieve its current code in the runtime memory and modify it by monkey patching, which dynamically modifies classes or modules during execution. In our implementation, we adhere to a minimalist design to minimize the influence of human priors. We implement the optimization module using a recursive function. In this module, LLM analyzes and makes a series of decisions, including reading and modifying its own code from runtime memory (self-awareness and self-modification), executing Python or Linux commands, and interacting with the environment to gather feedback. The agent then proceeds to the subsequent recursive depth and continues to optimize itself. It is worth noting that the optimization module may have already been modified by the time the recursion occurs, potentially enhancing its optimization capabilities.
@@
To validate the effectiveness of Gödel Agent, we conduct experiments on multiple domains including coding, science, math, and reasoning. Our experimental results demonstrate that Gödel Agent achieves significant performance gain across various tasks, surpassing various widely-used agents that require human design. The same implementation of Gödel Agent can easily adapt to different tasks by only specifying the environment description and feedback mechanism. Additionally, the case study of the optimization progress reveals that Gödel Agent can provide novel insights into agent design. We also investigate the impact of the initial policy for improvement on subsequent outcomes, finding that a good start can significantly accelerate convergence during optimization.
@@
  • We propose the first self-referential agent framework, Gödel Agent, based on LLMs. It autonomously engages in self-awareness, self-modification, and recursive self-improvement across any task, reducing the need for manual agent design and offering higher flexibility and freedom.
  • We implement Gödel Agent framework using the monkey patching method. Our experiments show that Gödel Agent outperforms manually designed agents and surpasses its earlier versions on several foundational tasks, demonstrating effective self-improvement.
  • We analyze Gödel Agent ’s optimization process, including its self-referential capabilities and the resulting agentic system, aiming to deepen our understanding of both LLMs and agentic systems.
@@
2 Method
@@
Algorithm 1 Recursive Self-Improvement of Gödel Agent
2:Output: Optimized policy $\pi$ and Gödel Agent $s$
[…]
5$\triangleright$ Compute the initial performance.
6:$r\leftarrow U(\mathcal{E},\pi_{0})$
7$\triangleright$ Perform recursive self-improvement.
8:$\pi,s\leftarrow\texttt{SELF\_IMPROVE}(\pi,s,r,g)$
9:return $\pi,s$
10:
11$\triangleright$ Initial code of self-referential learning.
12:function SELF_IMPROVE($\mathcal{E},\pi,s,r,g$)
13:     $\triangleright$ Obtain action sequence.
14:     $a_{1},\ldots,a_{n}\leftarrow f_{0}(\pi,s,r,g)$
15:     for $a_{i}$ in $a_{1},\ldots,a_{n}$ do
16:         $\pi,s,r\leftarrow\texttt{EXECUTE}(\mathcal{E},\pi,s,r,a_{i})$
17:     end for
18:     return $\pi,s$
19:end function
20:
21$\triangleright$ Initial action execution function.
22:function EXECUTE($\mathcal{E},\pi,s,r,a$)
23:     switch $a.\texttt{name}$
24:      case self_state:
[…]
26:      case interact:
27:       $r\leftarrow U(\mathcal{E},\pi)$
28:      case self_update:
29:       $\pi,s\leftarrow a.\texttt{code}$
30:      case continue_improve:
31:       $\triangleright$ Recursively invoke self-improvement.
32:       $\pi,s\leftarrow\texttt{SELF\_IMPROVE}(\mathcal{E},\pi,s,r,g)$
33:     return $\pi,s,r$
34:end function
@@
The agent code is updated to $(\pi_{t+1},I_{t+1})$ after the current execution of $(\pi_{t},I_{t})$ is finished. Both the agent algorithm $(\pi,I)$ and the action set $\mathcal{A}$ are not static and can be expanded and modified by the agent itself at the training time. Algorithm 1 illustrates the described algorithm for the Gödel Agent. Each recursive call enables the agent to refine its performance and become progressively more efficient.
3 Gödel Agent Initialization
@@ 3.1 Implementation Details
Self-Awareness via Runtime Memory Inspection  Our Gödel Agent achieves self-awareness by inspecting runtime memory, particularly local and global variables in Python. This capability allows the agent to extract and interpret the variables, functions, and classes that constitute both the environment and the agent itself, according to the modular structure of the system. By introspecting these elements, the agent gains an understanding of its own operational state and can adapt accordingly.
@@
Environmental Interaction  To assess performance and gather feedback, Gödel Agent is equipped with interfaces for interacting with its environment. Each task provides tailored environmental interfaces, enabling it to evaluate its performance and adjust its strategies accordingly. This interaction is a crucial part of the feedback loop in the recursive improvement process.
Recursive Improvement Mechanism  At each time step, Gödel Agent determines the sequence of operations to execute, which includes reasoning, decision-making, and action execution. After completing the operations, Gödel Agent evaluates whether its logic has improved and decides whether to proceed to the next recursive iteration. Over successive iterations, Gödel Agent’s logic evolves, with each step potentially improving its decision-making capacity.
Goal Prompt and Task Handling  The goal prompt informs Gödel Agent that it possesses the necessary privileges to enhance its logic and introduces the available tools for improvement. As shown in Appendix A, this prompt encourages Gödel Agent to fully explore its potential and leverage the tools for self-optimization. To ensure effectiveness across diverse tasks, we provide Gödel Agent with an initial policy, where it will start to explore different policies to analyze its efficiency in optimizing performance.
3.2 Additional Designs to Support Gödel Agent’s Optimization
@@
Error Handling Mechanism  Errors during execution can lead to unexpected terminations of the agent process. To mitigate this, we implement a robust error recovery mechanism. If an operation results in an error, Gödel Agent halts the current sequence and moves on to the next time step, carrying forward the error information to improve future decisions.
@@
Although these additional tools are not strictly necessary for self-improvement, their inclusion accelerates the convergence of Gödel Agent ’s recursive optimization process. We conducted ablation studies to assess the effectiveness of these tools, as discussed in Section 5.1.
4 Experiments
4.1 Baseline Methods
To establish a comprehensive baseline, we select both fixed hand-designed methods and a representative automated agent design technique. Our hand-designed methods are well-known approaches that focus on enhancing reasoning and problem-solving capabilities. These include: 1) Chain-of-Thought (CoT) (Wei et al., 2022) that encourages agents to articulate their reasoning processes step-by-step before providing an answer. 2) Self-Consistency with Chain-of-Thought (CoT-SC) (Wang et al., 2023b) that generates multiple solution paths using the CoT framework and selects the most consistent answer. 3) Self-Refine (Madaan et al., 2024) that involves agents assessing their own outputs and correcting mistakes in subsequent attempts. 4) LLM-Debate (Du et al., 2023) that allows different LLMs to engage in a debate, offering diverse viewpoints. 5) Step-back Abstraction (Zheng et al., 2024) that prompts agents to initially focus on fundamental principles before diving into task details. 6) Quality-Diversity (QD) (Lu et al., 2024) that generates diverse solutions and combines them. 7) Role Assignment (Xu et al., 2023) that assigns specific roles to LLMs to enhance their ability to generate better solutions by leveraging different perspectives. Given the limitations of fixed algorithms in handling dynamic scenarios, we select 8) Meta Agent Search (Hu et al., 2024), the latest state-of-the-art method for automated agent design, as our main comparison point.
4.2 Experimental Settings
Following the setup of Hu et al. (2024), we evaluate Gödel Agent’s self-improvement capabilities across four well-known benchmarks. The benchmarks are as follows: 1) DROP (Dua et al., 2019) for reading comprehension. 2) MGSM (Shi et al., 2022) for testing mathematical skills in a multilingual context. 3) MMLU (Hendrycks et al., 2021) for evaluating multi-task problem-solving abilities. 4) GPQA (Rein et al., 2023) for tackling challenging graduate-level science questions.
Given the complexity of the tasks and the need for advanced reasoning and understanding, the improvement cycle of Gödel Agent is driven by GPT-4o. In the main experiment, we implement two different settings: 1) To make a fair comparison with baseline methods, we forbid Gödel Agent to change the API of the LLM used to perform the tasks (by default GPT-3.5) and use a closed-book approach with no access to the Internet, and 2) To explore the upper bound of Gödel Agent’s capabilities, we remove all constraints. Chain of Thought is applied as the initial policy for all tasks, given its simplicity and versatility. In addition, as shown in Section 5.3, we also analyze the performance of Gödel Agent when using other algorithms as the initial policies.
We perform 6 independent self-improvement cycles for each task. Each cycle represents a complete self-improvement process, where Gödel Agent iteratively modifies its logic to enhance performance. Further details regarding the experimental setup and additional results can be found in Appendix B.
Table 1: Results of three paradigms of agents on different tasks. The highest value is highlighted in bold, and the second-highest value is underlined. Gödel-base is the constrained version of Gödel Agent, allowing for fair comparisons with other baselines. Gödel-free represents the standard implementation without any constraints, whose results are italicized. We report the test accuracy and the 95% bootstrap confidence interval on test sets [^3] The results of baseline models are refer to Hu et al. (2024)..

### v3

# Gödel Agent: A Self-Referential Agent Framework for Recursively Self-Improvement

@@ Abstract
The rapid advancement of large language models (LLMs) has significantly enhanced the capabilities of agents across various tasks. However, existing agentic systems, whether based on fixed pipeline algorithms or pre-defined meta-learning frameworks, cannot search the whole agent design space due to the restriction of human-designed components, and thus might miss the more optimal agent design. In this paper, we introduce Gödel Agent, a self-evolving framework inspired by the Gödel machine, enabling agents to recursively improve themselves without relying on predefined routines or fixed optimization algorithms. Gödel Agent leverages LLMs to dynamically modify its own logic and behavior, guided solely by high-level objectives through prompting. Experimental results on multiple domains demonstrate that implementation of Gödel Agent can achieve continuous self-improvement, surpassing manually crafted agents in performance, efficiency, and generalizability. Our code can be found at https://github.com/Arvid-pku/Godel_Agent.
Gödel Agent: A Self-Referential Agent Framework for Recursively Self-Improvement
Xunjian Yin111 Work completed during internship at UCSB. , Xinyi Wang , Liangming Pan , Li Lin Xiaojun Wan , William Yang Wang Peking University  University of California, Santa Barbara   University of Arizona {xjyin,efsotr_l,wanxiaojun}@pku.edu.cnxinyi_wang@ucsb.edu liangmingpan@arizona.eduwilliam@cs.ucsb.edu
@@ 1 Introduction
As large language models (LLMs)  (OpenAI et al., 2024; Dubey et al., 2024) demonstrate increasingly strong reasoning and planning capabilities, LLM-driven agentic systems have achieved remarkable performance in a wide range of tasks (Wang et al., 2024a). Substantial effort has been invested in manually designing sophisticated agentic systems using human priors in different application areas. Recently, there has been a significant interest in creating self-evolving agents, that not only greatly reduce human labor but also produce better solutions. Given that human effort can only cover a small search space of agent design, it is reasonable to expect that a self-evolving agent with the freedom to explore the full design space has the potential to produce a more optimal solution.
Figure 1: Modular demonstration of Gödel Agent. Compared with traditional agents, its sensor and executor can read and write all of its own code.
@@
There is a large body of work proposing agents capable of self-refinement. Some agents are designed to iterate over a fixed routine consisting of a list of fixed modules, while some of the modules are capable of taking self- or environment feedback to refine their actions  (Chen et al., 2023b; Qu et al., 2024a; Yao et al., 2023). This type of agent, referred to as Hand-Designed Agent, is depicted as having the lowest degree of freedom in Figure 2. More automated agents have been designed to be able to update their routines or modules in some pre-defined meta-learning routine, for example, natural language gradients (Zhou et al., 2024), meta agent (Hu et al., 2024), or creating and collecting demonstrations  (Khattab et al., 2023). This type of agent, known as Meta-Learning Optimized Agents, is depicted as having the middle degree of freedom in Figure 2. However, there are inevitably some human priors involved in these agent designs that cannot be improved during the inference time.
@@
Figure 2: Comparison of three agent paradigms. Hand-designed agents rely on human expertise which are limited in scope and labor-intensive. Meta-learning optimized agents are constrained by a fixed meta-learning algorithm, restricting their search space and optimization potential. In contrast, self-referential agent (Gödel Agent) can recursively improve itself without any limitation. Its optimization capabilities are constantly being enhanced by itself. Consequently, in return, it can continue to optimize itself better.
In this paper, we propose Gödel Agent to eliminate the human design prior, which is an automated LLM agent that can freely decide its own routine, modules, and even the way to update them. It is inspired by the self-referential Gödel machine (Schmidhuber, 2003), which was proven to be able to find the global optimal solutions. Self-reference means the property of a system that can analyze and modify its own code, including the parts responsible for the analysis and modification processes (Astrachan, 1994). Therefore, it can achieve what’s known as ”recursive self-improvement”, where it iteratively updates itself to become more efficient and effective at achieving its predefined goals. In this case, as shown in Figure 1, Gödel Agent can analyze and modify its own code, including the code for analyzing and modifying itself, and thus can search the full agent design space, which is depicted as having the highest degree of freedom in Figure 2. Gödel Agent can theoretically make increasingly better modifications over time through recursively self-update (Wang, 2018).
@@
In this paper, we choose to implement it by letting it manipulate its own runtime memory, i.e., the agent is able to retrieve its current code in the runtime memory and modify it by monkey patching (Bimal, 2012), which dynamically modifies classes or modules during execution. To allow it to update the logic of the running main function, unlike the loop-iterative approach of traditional agents, we implement the main function as a recursive function. In this function, LLM analyzes and makes a series of decisions, including reading and modifying its own code from runtime memory (self-awareness***In this paper, self-awareness means that the agent can introspect and read its own code and files, not to imply any philosophical sense of consciousness or awareness. and self-modification), and interacting with the environment to gather feedback. The agent then proceeds to the subsequent recursive depth and continues to optimize itself.
@@
To validate the effectiveness of Gödel Agent, we conduct experiments on multiple domains including coding, science, math, and reasoning. Our results demonstrate that Gödel Agent achieves significant performance gain across various tasks, surpassing various widely-used agents that require human design. The same implementation of Gödel Agent can easily adapt to different tasks by only specifying the environment description and feedback mechanism. Additionally, the case study of the optimization progress reveals that Gödel Agent can provide novel insights into agent design.
@@
  • We propose the first fully self-referential agent framework, Gödel Agent, and implement it using monkey patching. It autonomously engages in self-awareness, self-modification, and recursive self-improvement.
  • Experiments shows that Gödel Agent is superior to the previous agent frameworks in terms of performance, flexibility, cost, and potential.
  • We analyze Gödel Agent ’s optimization process, including its self-referential abilities and the optimized agentic systems, aiming to deepen our understanding of both LLMs and agents.
@@
2 Related Work
Hand-Designed Agent Systems  Researchers have designed numerous agent systems tailored to various tasks based on predefined heuristics and prior knowledge. These systems often employ techniques such as prompt engineering (Chen et al., 2023a; Schulhoff et al., 2024), chain-of-thought reasoning and planning (Wei et al., 2022; Yao et al., 2022), as well as reflection (Shinn et al., 2024; Madaan et al., 2024), code generation (Wang et al., 2023a; Vemprala et al., 2024), tool use (Nakano et al., 2021; Qu et al., 2024a), retrieval-augmented generation (Lewis et al., 2020; Zhang et al., 2024b), and multi-agent collaboration (Xu et al., 2023; Wu et al., 2023; Qian et al., 2023; Hong et al., 2023). Once crafted by human designers, these systems remain static and do not adapt or evolve over time.
Meta-Learning Optimized Agent Systems  Some researchers have explored methods for enhancing agents through fixed learning algorithms (Zhou et al., 2024; Hu et al., 2024). For example, certain frameworks store an agent’s successful or failed strategies in memory based on environmental feedback (Liu et al., 2023; Hu et al., 2023; Qian et al., 2024), while others automatically optimize agent prompts (Khattab et al., 2023; Zhang et al., 2024a; Khattab et al., 2023). Some studies focus on designing prompts that enable agents to autonomously refine specific functions (Zhang et al., ). However, these meta-algorithms are also designed manually and remain unchanged once deployed, limiting the agents’ ability.
Recursive Self-Improvement  The concept of recursive self-improvement has a long history (Good, 1966; Schmidhuber, 1987). Gödel machine (Schmidhuber, 2003) introduced the notion of a proof searcher that executes a self-modification, thereby enabling the machine to enhance itself. In the early days, there were also some discussions of self-improving agents that were not based on LLM (Hall, 2007; Steunebrink and Schmidhuber, 2012). More recently, Zelikman et al. (2023) applied recursive self-improvement to code generation, where the target of improvement was the optimizer itself. Some work (Havrilla et al., 2024; Qu et al., 2024b; Kumar et al., 2024) also explores recursive self-improvement by fine-tuning models to introspect and correct previous mistakes. Gödel Agent represents the first self-referential agent based on LLM. This approach is more flexible, removing human-designed constraints.
3 Self-Referential Gödel Agent
Algorithm 1 Recursive Self-Improvement of Gödel Agent
2:Output: Optimized policy $\pi$ and Gödel Agent $s$
[…]
5$\triangleright$ Compute the initial performance.
6:$r\leftarrow U(\mathcal{E},\pi_{0})$
7$\triangleright$ Perform recursive self-improvement.
8:$\pi,s\leftarrow\texttt{SELF\_IMPROVE}(\pi,s,r,g)$
9:return $\pi,s$
10$\triangleright$ Initial code of self-referential learning.
11:function SELF_IMPROVE($\mathcal{E},\pi,s,r,g$)
12:     $\triangleright$ Obtain action sequence.
13:     $a_{1},\ldots,a_{n}\leftarrow f_{0}(\pi,s,r,g)$
14:     for $a_{i}$ in $a_{1},\ldots,a_{n}$ do
15:         $\pi,s,r\leftarrow\texttt{EXECUTE}(\mathcal{E},\pi,s,r,a_{i})$
16:     end for
17:     return $\pi,s$
18:end function
19$\triangleright$ Initial action execution function.
20:function EXECUTE($\mathcal{E},\pi,s,r,a$)
21:     switch $a.\texttt{name}$
22:      case self_state:
[…]
24:      case interact:
25:       $r\leftarrow U(\mathcal{E},\pi)$
26:      case self_update:
27:       $\pi,s\leftarrow a.\texttt{code}$
28:      case continue_improve:
29:       $\triangleright$ Recursively invoke self-improvement.
30:       $\pi,s\leftarrow\texttt{SELF\_IMPROVE}(\mathcal{E},\pi,s,r,g)$
31:     return $\pi,s,r$
32:end function
@@ 3 Self-Referential Gödel Agent
The agent code is updated to $(\pi_{t+1},I_{t+1})$ after the current execution of $(\pi_{t},I_{t})$ is finished. Both the agent algorithm $(\pi,I)$ and the action set $\mathcal{A}$ are not static and can be expanded and modified by the agent itself at the training time. Algorithm 1 illustrates the described algorithm for the Gödel Agent. Each recursive call enables the agent to refine its logic and become progressively more efficient.
4 Gödel Agent Implementation
@@ 4.1 Implementation Details
Self-Awareness via Runtime Memory Inspection  Gödel Agent achieves self-awareness by inspecting runtime memory, particularly local and global variables in Python. This capability allows the agent to extract and interpret the variables, functions, and classes that constitute both the environment and the agent itself, according to the modular structure of the system. By introspecting these elements, the agent gains an understanding of its own operational state and can adapt accordingly.
@@
Environmental Interaction  To assess performance and gather feedback, Gödel Agent is equipped with interfaces for interacting with its environment. Each task provides tailored environmental interfaces, enabling it to evaluate its performance and adjust its strategies accordingly. In practical implementations, a validation set can be used to provide feedback.
Recursive Improvement Mechanism  At each time step, Gödel Agent determines the sequence of operations to execute, which includes reasoning, decision-making, and action execution. After completing the operations, Gödel Agent evaluates whether its logic has improved and decides whether to proceed to the next recursive iteration. Over the next iteration, the entire new logic will be applied.
Goal Prompt and Task Handling  The goal prompt informs Gödel Agent that it possesses the necessary privileges to enhance its logic and introduces available tools. As shown in Appendix A, the prompt encourages Gödel Agent to fully explore its potential and utilize tools for self-optimization. To ensure effectiveness across diverse tasks, we provide Gödel Agent with an initial policy, where it will start to explore different policies.
4.2 Additional Designs
@@
Error Handling Mechanism  Errors during execution can lead to unexpected terminations of the process. To mitigate this, we implement a robust error recovery mechanism. If an operation results in an error, Gödel Agent halts the current sequence and moves on to the next time step, carrying forward the error information to help future decisions.
@@
Although these additional tools are not strictly necessary for self-improvement, their inclusion accelerates the convergence of Gödel Agent ’s recursive optimization process. We conduct ablation studies to assess the effectiveness of these tools, as discussed in Section 6.1.
@@
Table 1: Results of three paradigms of agents on different tasks. The highest value is highlighted in bold, and the second-highest value is underlined. Gödel-base is the constrained version of Gödel Agent, allowing for fair comparisons with other baselines. Gödel-free represents the standard implementation without any constraints, whose results are italicized. We report the test accuracy and the 95% bootstrap confidence interval on test setsThe results of baseline models are refer to Hu et al. (2024)..
5 Experiments
5.1 Baseline Methods
To establish a comprehensive baseline, we select both hand-designed methods and automated agent design techniques. Hand-designed methods are well-known approaches that include: 1) Chain-of-Thought (CoT) (Wei et al., 2022) that encourages agents to reason step-by-step before providing an answer. 2) Self-Consistency with CoT (CoT-SC) (Wang et al., 2023b) that generates multiple solution paths using CoT and selects the most consistent answer. 3) Self-Refine (Madaan et al., 2024) that involves agents assessing their outputs and correcting mistakes in subsequent attempts. 4) LLM-Debate (Du et al., 2023) that allows different LLMs to engage in a debate, offering diverse viewpoints. 5) Step-back Abstraction (Zheng et al., 2024) that prompts agents to initially focus on fundamental principles before diving into task details. 6) Quality-Diversity (Lu et al., 2024) that generates diverse solutions and combines them. 7) Role Assignment (Xu et al., 2023) that assigns specific roles to LLMs to generate better solutions by leveraging different perspectives. Given the limitations of fixed algorithms in handling dynamic scenarios, we select 8) Meta Agent Search (Hu et al., 2024), the latest state-of-the-art method for automated agent design, as our main comparison point.
5.2 Experimental Settings
Following the setup of Hu et al. (2024), we evaluate Gödel Agent’s self-improvement capabilities across four well-known benchmarks: 1) DROP (Dua et al., 2019) for reading comprehension. 2) MGSM (Shi et al., 2022) for testing mathematical skills in a multilingual context. 3) MMLU (Hendrycks et al., 2021) for evaluating multi-task problem-solving abilities. 4) GPQA (Rein et al., 2023) for tackling challenging graduate-level science questions.
Given its simplicity and versatility, we use CoT as the initial policy for all tasks. In addition, as shown in Section 6.3, we also analyze the performance of Gödel Agent when using other algorithms as the initial policies.
We perform 6 independent self-improvement cycles on the validation dataset for each task, with a maximum of 30 iterations per cycle. Each cycle represents a complete self-improvement process, where Gödel Agent iteratively modifies its logic to enhance performance. After obtaining the optimized agent, we test it on the test set. For fairness, we use GPT-3.5 for all the tests, whether for the baseline or Gödel Agent. Further details can be found in Appendix B.